OpenStudy (anonymous):

rewrie the expression as an equivlent expression that does not contain powers of trigonometric functions greater than 1: 10sin ^4x

OpenStudy (anonymous):

Half-Angle Formula: \[ \begin{array}{} \sin^2(\theta) &=& \frac{1-\cos(2\theta)}{2}\\ \cos^2(\theta) &=& \frac{1+\cos(2\theta)}{2} \end{array} \]